For the function below find the inverse.

a. 𝑓(π‘₯) = (π‘₯ βˆ’ 2) 5 + 3

b. 𝑔(π‘₯) = 4 π‘₯+2

c. β„Ž(π‘₯) = 4 βˆ’ √4π‘₯ 3 2

d. 𝑓(π‘₯) = 4π‘₯ 3 + 2


1
Expert's answer
2021-03-30T16:08:43-0400

a). "y=8x-10" . We get the following inverse: "x=\frac{1}{8}(y+10)"

b). "y=4x+2" . We obtain the inverse: "x=\frac{1}{4}(y-2)"

c). "y=4-6\sqrt{4}x". The inverse is: "x=\frac{1}{6\sqrt{4}}(4-y)"

d). "y=12x+2" . We get the inverse: "x=\frac{1}{12}(y-2)"